Harrison Clothiers' stock currently sells for $39 a share. It just paid a dividend of $2.25 a share (that is, D0 = 2.25). The dividend is expected to grow at a constant rate of 5% a year.

a. What stock price is expected 1 year from now? Round your answer to two decimal places. $

b. What is the required rate of return? Round your answers to two decimal places. %

Ezzell Corporation issued perpetual preferred stock with a 9% annual dividend. The stock currently yields 9%, and its par value is $100.

a. What is the stock's value? Round your answer to two decimal places. $

b. Suppose interest rates rise and pull the preferred stock's yield up to 12%. What would be its new market value? Round your answer to two decimal places. $
Answer #1

1. Price at year 0 (P0) =39
Growth =5%
a. Stock price one year from now =P0*(1+growth) =39*(1+5%) =40.95
b. Required Rate =Dividend in year 1/Price +growth =2.25*(1+5%)/39+5% =11.06%

2. a. Stock Value =Annual Dividend/Current yield =9%*100/9%=100
b. New Market Value =Annual Dividend/Current yield =9%*100/12%=75